UT 28 Jun 2021



Night Info

  • Sunset 19:37
  • Twilight ends 21:13
  • Moon rises 22:27
  • Handover to PETS 22:00hr
  • Twlight stars 03:33
  • Sunrise 05:09

Night Plan

  • Bino-MODS to start
  • SpecPht Std in twilight as early as possible - Feige 67
  • MODS Science IMBH optical Spectra - Rothberg
    • Keep observing from SpecStd until reconfigure at 22:00hr
    • Conditions: Seeing < 1.2", up to 1mag cloud extinction
    • Scripts: /home/lbto/Service_Observing/2021A_D_E/May_Jun_Jul_2021/IMBH/ - OBS-N
    • OT File: JWST_PreFlight_Observations_V6.xml
  • Backup in case conditions too poor:
    • RLQ/RQQ Optical Spectra - Rothberg - any conditions - bino-MODS - can quickly shift to this program if conditions poor during SpecStd or acquisition of primary science
      • Script Location: /home/lbto/Service_Observing/2021A_D_E/May_Jun_Jul_2021/Rothberg_BW_Backup/QSOs
      • OT: RLRQQSO_V2.xml
      • Any of the three targets - any amount of time - options provided in case certain areas are windy/cloudy, etc.
      • Scripts have 2x1800sec exposures - if one can squeeze extra photons in before handover, please do.
        • PG 1416-129
        • PG 1307+085
        • HB89 1635+119

Overview:

0:32 Issues bringing MODS up on ROBS. Error starting modsDisp on ROBS:

Starting the modsDisp session...
  Launching MODS1R ds9 display, please wait...
Traceback (most recent call last):
  File "/lbt/lbto/mods/bin/modsDisp", line 375, in <module>
    md = start_ds9(ds9ID, timeout=15)
  File "/lbt/lbto/mods/bin/modsDisp", line 228, in start_ds9
    raise RuntimeError(f"failed to detect any ds9 window "
RuntimeError: failed to detect any ds9 window with title 'MODS1R-1624839244-3115673'

0:35 MODS brought up on OBS2. Awake and simSnap taken. Placed in observing mods for test preset.

Yang and Olga looked into the MODS Disp issues and it was working ok. I started it up and it was working as well. I had tried multiple times with the flags and was unable to start it earlier but cals were already in progress. Unsure why this occurred.

0:45 Taking Cals for IMBH J1201 and BW QSO backup observation:
Dual grating Slitless flats mods[1/2]b.20210628.20210627.0003-12, mods[1/2]r.20210628.20210627.0003-7
Dual grating Arcs mods[1/2]b.20210628.20210627.0013-15, mods[1/2]r.20210628.20210627.0008-10
Dual grating 0.6" flats mods[1/2]b.20210628.20210627.0016-22, mods[1/2]r.20210628.20210627.0011-16
Dual grating 5" flats mods[1/2]b.20210628.20210627.0023-27, mods[1/2]r.20210628.20210627.0017-19
8Kx3K biases mods[1/2]b.20210628.20210627.0028-37, mods[1/2]r.20210628.20210627.0020-29

2:30 Conditions overhead are clearing with some light cloud hanging around. There are thicker clouds to the south that Josh is monitoring. Humidity 36%, winds 4m/s from the south west.

2:42 Josh is opening us up.

2:50 and we are doing pointing checks now. Light clouds on the summit now but there is a cloud mass on the sattelite approaching the summit.

3:10 Sending acq preset to Feige 67. Some scattered light cloud, seeing 0.9-1.05" on the guiders.
3:20 Refined the alignment and waiting a few minutes for sky to come down a bit before starting specphot science. Seeing seems to be blowing up and is variable, 1.2- 1.8"

3:28 Starting science on Feige 67 (perhaps a bit early but will see and add exposures if needed). Seeing has clown up to 1.8-2.3". DIMM reporting 2.7-2.9"
mods[1/2]b.20210628.20210627.0038-41, mods[1/2]r.20210628.20210627.0033-36

3:46 Seeing is variable but calming down....now 1-1.3". Moving to primary science J1201.

3:53 Aborting seeing 2.13 at this lower elevation. Sending preset to PG1416-129. Seeing 1.45" on teh DIMM, 1.4-1.7" on the guiders. I missed the 120sec slit image in acquisition scripts when generating should be adjusted to 20sec for next time. 120sec is also a bit overkill for acquisition times on these bright objects.

4:08 Starting science on PG1416-129, seeing 1.46" on the DIMM, 1.6-1.7" on the guiders. Mostly clear
mods[1/2]b.20210628.20210627.0042, mods[1/2]r.20210628.20210627.0040

4:24 Seeing 3.6" on the DIMM and NaN on the guiders fora bit. Boucing around 2-3.5"

4:44 First exp complete. Handing over to PEPSI reconfig. Putting MODS 2 sleep. Hatch closed.

5:14 Reconfig issues so late start. First PEPSI target V1674 Her. Josh is sending pointing.

5:20 Going to a secondary pointing and collimation star. Seeing 1.7-2" on the DIMM and guiders

5:27 Bin 4 guiding on V1674 Her Mag 12.67, CD 3, CD 5, 300 fiber.

5:30 strating exposures. 2.25" on the DIMM, 2.5" on teh guiders. Exposure time on V1674 Her is upped from 15 min to 20 minutes

5:50 Sending preset to Kelt-9
5:53 starting exposure on Kelt 9 CD3 and CD5 200 fiber FPE time series with 5min exposures. Seeing 3.3-5" on the Dimm, 3.6-4.6" on the guiders.

Seeing ~3" and up to 4.5" for most of the second half Improving for the last hour of the night to 1.5"

3:35 12 degree twilight

-- BarryRothberg - 25 Jun 2021
Topic revision: r3 - 28 Jun 2021, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback